OpenStudy (anonymous):

solve! x^2 - 12x + 36 = 25 Will give medal!

OpenStudy (anonymous):

@wio ? can you help with this one also?

OpenStudy (anonymous):

Okay, so... before I said \[ x^2+bx = (x+b/2)^2-(b/2)^2 \]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

For \(x^2-12x\), we have \(b=-12\).

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Can you substitute that value for \(b\) into\[ (x+b/2)^2-(b/2)^2 \]?

OpenStudy (anonymous):

yeah so it would be x + -12/2)^2 - (12/2)^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can you simplify it a bit more?

OpenStudy (anonymous):

(x -6)^2 - (6)^2

OpenStudy (anonymous):

Good, also you can simplify the \(6^2\) as well, we just don't want to foil out \((x-6)^2\).

OpenStudy (anonymous):

so (x - 6)(x-6) ?

OpenStudy (anonymous):

No, don't foil that part out.

OpenStudy (anonymous):

My point is we can simplify it to: \[ (x-6)^2-36 \]

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Now, we can say: \[ x^2-12x = (x-6)^2-36 \]

OpenStudy (anonymous):

Can you plug that into the red part here?\[ \color{red}{x^2 - 12x} + 36 = 25 \]

OpenStudy (anonymous):

I'm kinda confused on what you mean...?

OpenStudy (anonymous):

Since we have: \[ x^2-12x = (x-6)^2-36 \]we can swap the two of them whenever we want. That means we can swap \(x^2-12x\), in the expression \[ \color{red}{x^2-12x}+36=25 \]with \((x-6)^2-36\) to get: \[ [\color{red}{(x-6)^2-36}]+36 = 25 \]

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

Now, the \(-36+36\) cancel out, and we are left with: \[ (x-6)^2=25 \]

OpenStudy (anonymous):

okay,..

OpenStudy (anonymous):

When we take the square root of both sides, we end up with: \[ |x-6|=5 \]

OpenStudy (anonymous):

Are you familiar with the absolute value bars?

OpenStudy (anonymous):

wooaahhh... Is that absolute value? or did you mean parenthesis

OpenStudy (anonymous):

Yes, it is the absolute value.

OpenStudy (anonymous):

wouldnt you just square root the other side?

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

Here is the truth: \[ \sqrt{(\ldots)^2} = |\ldots| \]

OpenStudy (anonymous):

Hmmm. But this is for homework and I dont recall us going over the absolute values... but ok

OpenStudy (anonymous):

So \[ \sqrt{25} = |5| = 5 \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Well, sometimes they forget to mention the absolute value, but it is there.

OpenStudy (anonymous):

So do you remember absolute value at all, or do I need to give you a reminder?

OpenStudy (anonymous):

no i remember but now we just add 6?

OpenStudy (anonymous):

so x = 11 ?

OpenStudy (anonymous):

The absolute value splits it into two equations.

OpenStudy (anonymous):

The first equation is the positive one: \[ x-6 = 5 \]The second equation is the negative one: \[ -(x-6)=5 \]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

What answers did you get?